### 06 Nov 2024 v1

# A Hierarchy of Spectral Gap Certificates for Frustration-Free Spin Systems

Abstract
Estimating spectral gaps of quantum many-body Hamiltonians is a highly challenging computational task, even under assumptions of locality and translation-invariance. Yet, the quest for rigorous gap certificates is motivated by their broad applicability, ranging from many-body physics to quantum computing and classical sampling techniques. Here we present a general method for obtaining lower bounds on the spectral gap of frustration-free quantum Hamiltonians in the thermodynamic limit. We formulate the gap certification problem as a hierarchy of optimization problems (semidefinite programs) in which the certificate—a proof of a lower bound on the gap—is improved with increasing levels. Our approach encompasses existing finite-size methods, such as Knabe’s bound and its subsequent improvements, as those appear as particular possible solutions in our optimization, which is thus guaranteed to either match or surpass them. We demonstrate the power of the method on one-dimensional spin-chain models where we observe an improvement by several orders of magnitude over existing finite size criteria in both the accuracy of the lower bound on the gap, as well as the range of parameters in which a gap is detected.
I Introduction
A central question in many-body quantum physics is to determine whether a system is gapped or gapless in the thermodynamic limit. Indeed, the existence of a gap has important consequences for the physical properties of the low energy sector. For instance, it controls correlations and entanglement in the ground state, both in 1D [1, 2] and in some 2D systems [3]. Moreover, in 1D it directly relates to a notion of bounded complexity [4]. Estimating the gap is also crucial to several tasks, including the classification of quantum phases of matter [5, 6], adiabatic state preparation algorithms [7, 8] and the identification of efficient Monte Carlo sampling methods [9, 10].
Proving the existence of a spectral gap is a highly non-trivial mathematical challenge. Even for reasonable local Hamiltonians the task can be undecidable [11, 12, 13]. Several techniques to lower bound the spectral gap have been developed for the better-behaved class of frustration-free Hamiltonians, i.e. systems in which the ground state minimizes each local term individually. The most notable lower bounds are based on the “martingale” method [14] and finite size criteria [15, 16]. More recently, methods inspired by the numerical bootstrap approach to conformal field theory have been proposed to determine spectral properties of simple quantum Hamiltonians as well as quantum many body systems [17, 18, 19].
The martingale method is a powerful and widely applicable analytical technique. In particular, using this approach, the question of the existence of a gap has been solved for a wide class of for 1D frustration-free models [16, 14]. Such methods, powerful as they are, fall short in that they do not always allow to get an actual lower bound for the gap, and in cases in which a number can be obtained, it is typically not a particularly tight lower bound.
Finite-size criteria, on the other hand, provide more realistic lower bounds on the spectral gap, with practical computational requirements. Such methods establish a relation between the gap of the infinite system and a quantity pertaining to a finite-size system (with the same interaction). This approach was pioneered by Knabe who was able to provide a lower bound on the gap of the AKLT chain [15]. Its subsequent improvements and generalizations [20, 21, 22, 23, 24] have allowed to prove the gap in a wide range of systems relevant to both many-body physics [25, 26, 27, 28, 29, 30, 31] and quantum information [32, 33]. Despite recent advances [34, 35, 36], crucial questions remain open, particularly for systems in spatial dimensions larger than one.
In this work we put forward a systematic way to exhaust the search for gap certificates. Generalizing the the core idea underlying both the martingale approach and finite-size criteria, we construct a hierarchy of relaxations of the gap-estimation problem. Each level $n$ in the hierarchy is a finite $n$-body optimization problem, the solution to which is a lower bound on the infinite-system gap. As the level is increased the quality of the lower bound is systematically improved. Crucially, we are able to prove that our method always matches or outperforms all finite-size criteria we know of, including the Knabe [15] and the Gosset–Mozgunov [20] bounds, as those appear as possible (but not necessarily optimal) solutions in our optimization problem.
We demonstrate the power of our method on several paradigmatic examples of 1D frustration-free models. Our method leads to the most accurate lower bound for the spin-1 AKLT chain do date, and detects the gap in parameter regimes far beyond those where other finite-size criteria can detect. While the presentation in the paper is specialized to 1D systems, the approach can be straightforwardly generalized to higher-dimensional lattices.
To put our results into context, we find it worthwhile to pause and briefly discuss the basic idea behind our method—and how it generalizes existing gap certification methods—on a slightly more technical level. The well known idea underlying all gap certification approaches is the fact that a frustration-free Hamiltonian $H$ has a gap greater or equal to $\delta$ if the operator $H^{2}-\delta H$ is positive semidefinite (i.e. all its eigenvalues are non negative). We notice that all of the above mentioned gap certification methods can be regarded as finding different ways to decompose $H^{2}-\delta H$ as a sum of positive terms. Inspecting the different known methods, we notice that in order to attain better lower bounds it seems to be necessary to allow for more freedom in choosing the form of the terms in the decomposition. A good example of this is Gosset and Mozgunov’s [20] improvement over Knabe’s bound [15] which was obtained by introducing additional free parameters in what is otherwise a very similar proof. This observation leads us to formulate the gap estimation problem as an optimization problem where $\delta$ is maximized under the constraint that $H^{2}-\delta H$ can be written as a sum of positive terms. A first version of this idea has been proposed in Ref. [37]. To exhaust the search space in a systematic way we organize it as hierarchy of optimization problems: at each level of the hierarchy we optimize over all $n$-local positive decompositions of $H^{2}-\delta H$.
The rest of the paper is organized as follows: In Section II, we discuss the main technical ingredients of our method. In Section III, we show how previous finite-size criteria arise as special cases of the semidefinite program introduced in Section II. In Section IV, we present numerical results comparing lower bounds obtained using our method against those obtained with existing finite-size criteria.
II The method
We consider a spin chain comprised of particles defined on a $d-$dimensional Hilbert space $\mathbb{C}^{d}$ each. A translationally-invariant (TI) $k-$body Hamiltonian is defined as the operator
$H=\sum_{i=1}^{N}h_{i},$ (1)
[…]
Without loss of generality, we will shift the ground state energy to be 00, which, combined with frustration-freeness, implies both $H\succeq 0$ and $h_{i}\succeq 0$, where $\succeq$ is used to denote that the matrix is positive semidefinite.
The spectral gap is then the lowest non-zero eigenvalue of $H$ and we denote it by $\Delta_{N}^{\circ}$, where the $\circ$ superscript symbolizes periodic boundary conditions. Of particular relevance is the behavior of $\Delta_{N}^{\circ}$ as the system size increases. A model is defined to be gapped if the spectral gap does not close in the thermodynamic limit, that is $\Delta_{N\to\infty}^{\circ}=\Delta>0$.
We now make use of a well-known connection between the existence of a spectral gap and a quadratic operator inequality in terms of $H$: A frustration-free Hamiltonian has a gap greater or equal to $\delta$ if and only if the Hamiltonian satisfies the condition
$H^{2}-\delta H\succeq 0.$ (2)
The above condition enforces that $H$ has no eigenvalues in the interval $(0,\delta)$. Let us stress that all the most commonly-used methods to prove the gap of a system rely on establishing condition (2) for a given value of $\delta$ [15, 23, 14, 21, 36, 34, 35].
The condition in eq. 2 allows us to formulate the gap estimation problem as an optimization problem, which truns out to be a semidefinite program (SDP).
$\Delta_{N}^{\circ}$ $=\max\delta\,\,$
s.t. $\quad H^{2}-\delta H\succeq 0,$ (3)
under the condition that the ground state energy of $H$ is zero. The optimal value of this SDP is exactly the energy of the first excited state of $H$. Additionally, at the optimal solution, both the ground state and the first excited state space of $H$ lie in the ground state space of $H^{2}-\Delta H$.
In the general case, solving this SDP is as hard as exactly diagonalizing the Hamiltonian, since this problem imposes a positivity constraint on a matrix of size $(d^{N}\times d^{N})$. As $N$ increases, very quickly it becomes computationally intractable to solve. In the following, we show how to relax this problem. The resulting relaxation is a local, $n$-body, problem the solution to which is a lower bound on the gap of $N$ site periodic system for all $N\geq 2n$, and therefore holds for the infinite system as well.
II.1 Simplifications of the general SDP
In this subsection we explain how the gap problem in an infinite translationally invariant system can be solved locally. To do this we will first truncate the operator in eq. 2, replacing it with an $n$-local operator by omitting some positive terms from $H^{2}$. The positivity of the omitted terms will ensure that we still obtain a lower bound on the gap in the end. Second, we will formulate a finite size ($n$-body) problem which looks for a positive local term which, when summed over all translations, gives rise to the truncated $H^{2}-\delta H$ operator. Let us explain those two steps in detail (cf. fig. 1).
Figure 1: Illustration of the two relaxation steps. (a) Geometrical locality condition: The terms discarded from $H^{2}$ in eq. 4 are positive semi-definite since $h_{i}$ and $h_{j}$ have disjoint support. (b) Freedom in choosing a local generating term: Two local terms $g_{n}$ and $q_{n}$ generating the same global operator can differ as in eq. 10 as this difference cancels telescopically. Inside the $n=5$-site support of each local term (orange rectangle), the operators $Y_{n-1}$ are depicted as rectangles of different colors and the identity operator is depicted by a gray circle. When summed over all $i$, the terms colored identically cancel out as indicated by the crosses. The purple and gray terms are canceled by the preceding and the following terms in the sum respectively.
Geometric Locality condition. The first problem we need to address is that $H^{2}$ is not local: it consists of products of interactions terms $h_{i},h_{j}$ arbitrarily far apart. We therefore introduce a truncated operator $\left[H^{2}\right]_{n}$ which is $n$-local, by removing terms that act on spins farther than $n$ sites apart, namely
$\begin{split}\left[H^{2}\right]_{n}&=\sum_{|i-j|<n}h_{i}h_{j}\\ &=\sum_{i}\left(h_{i}^{2}+\sum_{i<j<i+n-k}\{h_{i},h_{j}\}\right).\end{split}$ (4)
If $n$ is chosen suitably (i.e. $n\geq 2k+1$), terms $h_{i}$ and $h_{j}$ with $j\geq i+n$ act on disjoint sets of spins and thus every discarded term is positive semidefinite (as a tensor product of positive terms). It follows that, if we define the following operator,
$Q_{n}(\delta)$ $:=\left[H^{2}\right]_{n}-\delta H\,$ (5)
the positivity of $Q_{n}(\delta)$ implies positivity of $H^{2}-\delta H$. Hence, we can prove condition (2) by showing $Q_{n}(\delta)\succeq 0$.
Positive semidefinite generating local term. Next, we will try to prove $Q_{n}(\delta)\succeq 0$ by looking for an $n$-body local term $q_{n}$ which is positive semidefinite such that $Q_{n}(\delta)$ can be written as
$Q_{n}(\delta)=\sum_{i}q_{n,i}\,\,,\quad q_{n,i}\succeq 0$ (6)
where the the $i$ subscript indicates the position of the first spin on which $q_{n}$ acts (the $n$ subscript simply denotes the $n$-locality, i.e. that $q_{n}$ is supported on $n$ neighboring sites). If for some value of $\delta$ $Q_{n}(\delta)$ is positive semidefinite, we can hope to find such a positive generating term because $Q_{n}(\delta)$ is itself $n$-local and translational invariant.
Note that if (6) holds for a given support $n$ and a global system size $N$, then the same decomposition will hold for any $N^{\prime}>N$. Hence, finding such a decomposition with a positive $q_{n}$ serves as a proof for the infinite system as well.
We will now proceed to show how the condition (6) can be expressed in terms of local operators, and thus arrive at a finite size relaxation of section II that is numerically tractable for any finite support $n$.
II.2 Primal SDP
After substituting the operator $H^{2}$ in the exact SDP section II with $[H^{2}]_{n}$, and imposing the condition that $Q_{n}(\delta)$ has a positive operator decomposition we arrive at the following problem
$\max_{\delta,q_{n}}\delta\,\,$
s.t. $\quad Q_{n}(\delta)=\sum_{i=1}^{N}q_{n,i},$ (7)
$\quad q_{n}\succeq 0\,.$ (8)
At this stage, the constraint eq. 7 still involves the exponentially-big operator $Q_{n}(\delta)$. We now argue that the condition (7) is equivalent to a local one, involving only operators with support on $n$ sites. First, notice that $Q_{n}(\delta)$ can be written in terms of a local generating term $g_{n}$ as $Q_{n}(\delta)=\sum_{i}g_{n,i}(\delta)$ where
$g_{n}(\delta):=\left(h_{1}^{2}+\sum\limits_{1<j<1+n-k}\{h_{1},h_{j}\}-\delta h_{1}\right)\,.$ (9)
The above term $g_{n}$ will in general not be positive semidefinite, even for a globally positive semidefinite operator $Q_{n}(\delta)$. Moreover the choice of local generating term is not unique as we can choose different ways to distribute the coefficients of the terms in $g_{n}$ as long as they sum up to the right amount in the global operator. For instance, the two local terms, $h_{i}=Z_{i}Z_{i+1}+X_{i}$ and $h^{\prime}_{i}=Z_{i}Z_{i+1}+\frac{1}{2}X_{i}+\frac{1}{2}X_{i+1}$, both give rise to the transverse field Ising model Hamiltonian when summed over all sites in the system. This freedom of choice can be completely characterised, which is done in Appendix A. In particular, we show that two generating local terms $q_{n}$ and $g_{n}$ sum up to the same global operator, if and only if they are related by
$q_{n}=g_{n}+\mathbbm{1}\otimes Y_{n-1}-Y_{n-1}\otimes\mathbbm{1}\,,$ (10)
where $Y_{n-1}$ can be any operator supported on $n-1$ sites. Condition (10) is chosen such that, when summed over $i$, the terms involving $Y_{n-1}$ cancel telescopically. It follows that the constraint (7) corresponds to looking for a suitable $Y_{n-1}$ that makes the local generating term $q_{n,i}$ positive semidefinite. In this way, we arrive at the final form of our SDP, which reads
$\Delta_{\mathrm{LTI}}(n)$ $=\max_{\delta,Y_{n-1}}\;\delta$ (11)
s.t. $\quad g_{n}(\delta)+\mathbbm{1}\otimes Y_{n-1}-Y_{n-1}\otimes\mathbbm{1}\succeq 0,$ (12)
where $g_{n}$ can be any local generating term for $Q_{n}(\delta)$, for example the one in eq. 9. Any two generating terms are related by some $Y_{n-1}$ terms as in eq. 10, and we anyway optimize over all possible choices of those terms.
We will refer to the problem we arrived at in eq. 11 as the locally translationally invariant (LTI) gap SDP. It is a semidefinite program which involves a positivity constraint on an $n$-body (i.e. $d^{n}\times d^{n}$ dimensional) matrix.
The LTI SDPs form a hierarchy of optimization problems, and the level of the hierarchy is labeled by the size $n$ of the finite system. As the size $n$ of the local operators increases, the solution $\Delta_{\mathrm{LTI}}(n)$ is guaranteed not to decrease, forming a sequence of lower bounds on the gap of the infinite system
$\Delta_{\mathrm{LTI}}(n)\leq\Delta_{\mathrm{LTI}}(n+1)\leq\ldots\leq\Delta_{N\to\infty}\,.$ (13)
This follows from the fact that the operator $q_{n+1}$ supported on $n+1$ sites embeds the solution $q_{n}$ obtained using $n$ sites, and they are related as
$q_{n+1}=a\cdot\mathbbm{1}\otimes q_{n}+b\cdot q_{n}\otimes\mathbbm{1}+c\cdot S,$ (14)
where $a,b,c\geq 0$ and $S\succeq 0$. Thus, the SDP provides more accurate lower bounds as $n$ increases, albeit at a higher computational cost.
II.3 Dual SDP
Any semidefinite program which is defined as a maximization has a dual formulation in terms of a minimization problem [38]. Importantly, the optimum of the dual serves as an upper bound on the primal optimal. When the two optima coincide, the problem exhibits so-called strong duality and one can equivalently solve either one of the problems. In this subsection we analyse the dual problems related to the spectral gap SDPs introduced in the previous subsections and show how they can provide quantitative information about the first excited state.
The dual program of the exact spectral gap SDP of section II is
[…]
s.t. $\rho\succeq 0$
[…]
This problem has strong duality, because the choice $\rho=\ket{\psi_{1}}\bra{\psi_{1}}/\Delta$, where |$\ket{\psi_{1}}$ denotes the first excited state of $H$; is feasible and attains the solution $\Delta$, which is the optimal solution of the primal. An alternate way to show that the optimal solution for the dual is directly related to the first excited state is highlighted in Appendix B.
This formulation of the gap problem provides a physical interpretation to the truncation step in eq. 4, where we disregarded far away terms $\{h_{i},h_{j}\}$ in $H^{2}$. As we see in eq. 15 the exact gap is given by the expectation value of $H^{2}$ in the first excited state. If that state exhibits fast decaying correlations then the contribution of the far away terms we neglected should not noticeably change the result.
Analogously to eq. 15, the dual of the TI-SDP problem has the form
[…]
s.t. $\rho_{n}\succeq 0$
[…]
where the optimization variable $\rho_{n}$ corresponds to the reduced marginals on $n$ parties. The constraint
[…]
is called local translation invariance (LTI) condition and arises as a dual to the constraint eq. 12. Moreover due to strong duality, the following relation is satisfied between the primal and dual optimal variables,
$\rho_{n}q_{n}=0.$ (18)
This relation which is known as complementarity slackness has several useful consequences. Among others, it provides us with the gradient of the solution to the SDP with respect to parameters in the problem data, without incurring further computational cost [39]. In our case the data is the Hamiltonian interaction term. We can then use the gradient to optimize the gap over all parent Hamiltonians of a given ground state. We elaborate on this in Appendix D.
[…]
III Connection to previous finite-size criteria
In this section, we compare our method to finite-size criteria-based techniques, which are the prevalent methods used to lower bound spectral gaps. For consistency with the formulation of the existing methods[15, 20], we restrict the discussion to nearest-neighbor projector Hamiltonians
$H=\sum_{i}h_{i},\quad h_{i}^{2}=h_{i}.$ (19)
Importantly, if the local term $h$ of a given frustration-free Hamiltonian is not a projector, it can always be modified into a projector Hamiltonian by adjusting the eigenvalues of the subspace orthogonal to the ground space of $h$. It follows from a simple argument that proving the gap of this second Hamiltonian implies a gap for the first as well [20](c.f. Section IV). We will now prove that the set of feasible local terms in the LTI-SDP eq. 11 includes the ones obtained from known finite size criteria. Thus, the value of LTI SDP achieves a tighter lower bound on the spectral gap.
III.1 Finite-size criteria
In the following discussion, we consider a nearest neighbor Hamiltonian comprised of local terms that are projectors, as in (19). We review the existing methods based on finite-size conditions, classifying them into two types - based on (i) finite system gaps, and (ii) local ground state projectors.
(Method 1) Based on finite system gaps. First formulated by Knabe [15], these criteria provide a way to relate the spectral gap of a finite open-boundary system, to the spectral gap in the thermodynamic limit. The open boundary Hamiltonian on a finite system of size $n$ is,
$h_{n,i}=\sum\limits_{j=0}^{n-1}h_{i+j},$ (20)
and $i$ denotes the site index where the support of the open boundary Hamiltonian begins. If $h_{n,i}$ has a spectral gap of $\epsilon_{n}$, then the lower bound to the gap in the thermodynamic limit is given as
$\Delta(n)=f(n)\left(\epsilon_{n}-g(n)\right).$ (21)
If $\Delta_{m}^{\circ}$ is the gap of periodic system on $m$ sites, then the Knabe bound [15] is
$\Delta_{m}^{\circ}\geq\Delta_{\mathrm{Knabe}}(n)=\frac{n-1}{n-2}\left(\epsilon_{n}-\frac{1}{n-1}\right)$ (22)
for $m>n>2$. Similarly, the Gosset–Mozgunov bound [20] is
$\Delta_{m}^{\circ}\geq\Delta_{\mathrm{GM}}(n)=\frac{5}{6}\frac{n^{2}+n}{n^{2}-4}\left(\epsilon_{n}-\frac{6}{n^{2}+n}\right)$ (23)
for $m>2n>4$. Comparing the bounds with eq. 21, the functions $f(n)$ and $g(n)$ satisfy the following properties
$\lim_{n\to\infty}f(n)\leq 1,\quad\lim_{n\to\infty}g(n)=0.$ (24)
such that
$\lim_{n\to\infty}\Delta_{\mathrm{Knabe},\mathrm{GM}}(n)\leq\lim_{n\to\infty}\epsilon_{n}.$ (25)
Moreover, for the case of gapless models, the function $g(n)$ provides an upper bound on spectral gap scaling with $n$ for the open boundary Hamiltonian $h_{n,i}$. The Gosset–Mozgunov bound improved this upper bound to $\mathcal{O}(n^{-2})$ from $\mathcal{O}(n^{-1})$ concluded from the Knabe bound. In a recent article, this $\mathcal{O}(n^{-2})$ scaling was generalized to arbitrary graphs and lattice dimensions [41].
(Method 2) Based on ground state projectors.
This method was introduced by Nachtergaele, Fannes, and Werner in Ref. [14, 16], and is also known as the ”martingale” criterion. Instead of using the gap of a finite size system, the presence of a gap is established by checking if a relation in terms of the ground space projectors is satisfied. The bound is obtained by first constructing a new Hamiltonian $\tilde{H}=\sum_{i}\tilde{h}_{\mathcal{I},\mathcal{I}+1}$ obtained by blocking $l$ sites into a single site, i.e.
$\{i,...,i+l-1\}$ $\to\mathcal{I}$ (26)
$\tilde{h}_{\mathcal{I},\mathcal{I}+1}=\sum\limits_{x=0}^{l+1}h_{i(l+1)-l+x}$ $\Rightarrow H\geq\frac{\tilde{H}}{2}.$ (27)
W.l.o.g, the blocked Hamiltonian $\tilde{h}$ can be converted into the corresponding projector Hamiltonian $K=\sum_{i}P_{\mathcal{I},\mathcal{I}+1}$, which satisfies the relation,
$\tilde{h}_{\mathcal{I},\mathcal{I}+1}\geq\epsilon_{2l}P_{\mathcal{I},\mathcal{I}+1},$ (28)
where $\epsilon_{2l}$ denotes the gap of $\tilde{h}$. To compute the martingale bound, we first state the inequality from Lemma 6.3(2) of [16],
$\{P_{\mathcal{I},\mathcal{I}+1},P_{\mathcal{I},\mathcal{I}+2}\}\succeq-\eta(P_{\mathcal{I},\mathcal{I}+1}+P_{\mathcal{I}+1,\mathcal{I}+2}),$ (29)
where
$\eta_{l}=||(\mathbb{I}-G_{1,\ldots,2l})(\mathbb{I}-G_{l+1,\ldots,3l})-(\mathbb{I}-G_{1,\ldots,3l})||,$ (30)
and $G_{i,i+1,...,i+m-1}$ denotes the projector onto the ground space of the original Hamiltonian $h$ on total $m$ sites from indices $i$ to $i+m-1$. Finally, using the above inequality, the bound on the gap of the original Hamiltonian is given as,
$\Delta_{M}(n=3l)=\frac{\epsilon_{2l}}{2}(1-2\eta_{l}),$ (31)
where $n$ denotes the number of sites in the original Hamiltonian. The lower bound of $\tilde{H}$ is $(1-2\eta)$ which is then rescaled by $\epsilon_{2l}/2$ to recover the bound for the original Hamiltonian $H$. Therefore, to prove the existence of a gap it is sufficient to find a value for $l$ that satisfies the condition $\eta_{l}<0.5$, which is useful in cases where analytical proofs are required [42, 32]. Given the form for the ground state projectors, it is possible to compute $\eta_{l}$ numerically by increasing the value of $l$ until the threshold of $\eta_{l}<0.5$ is achieved. The other possibility is to analytically bound $\eta_{l}$ as $\eta_{l}\leq\eta^{\prime}(l)$, where the function $\eta^{\prime}(l)$ becomes small enough for large $l$. This is problematic if the entire gap estimate (i.e. $\Delta_{M}$) is required, because, if the value of $l$ for which $\eta^{\prime}(l)<0.5$ is large, then it becomes harder to compute the prefactor $\epsilon_{2l}$ in eq. 31, as it requires the exact diagonalization of a $2l$-sites system.
III.2 Proof of inclusion
We will now prove that the LTI SDP provides more accurate and larger estimates for the thermodynamic limit gap than the finite size criteria bounds. More precisely, we obtain the relation,
$\Delta_{\mathrm{LTI}}(n)\geq\Delta_{\mathrm{Knabe,GM,M}}(n),$ (32)
meaning that the Knabe, Gosset–Mozgunov, and ”martingale” bounds (eqs. 21 and 31) calculated for a particular size $n$ of the truncation, yield weaker estimates, than the LTI SDP of the same size. Concretely, in SDP terms, we show that the finite size criteria bounds lie in the feasible set of the LTI SDP. The above relation relies on finding a positive semidefinite generating term $q_{n}(\delta)$ for the truncated operators $Q_{n}(\delta)$ corresponding to each of the bounds. Here we show the expressions for martingale bound,
$q_{n}(1-2\eta_{l}):=\eta_{l}(h_{i}+h_{i+1})+\sum_{x=1}^{n-2}\{h_{i},h_{i+x}\},$ (33)
which corresponds to the lower bound on the gap of the blocked Hamiltonian $K$ (as defined in Method 2 (section III.1)) and the expression for the Knabe bound is,
$q_{n}(\Delta_{\mathrm{Knabe}}(n))$ $:=\left(\frac{1-\epsilon_{n}}{n-2}\right)h_{n,i}$
$+\sum_{x=1}^{n-2}\frac{1}{n-x-1}\sum\limits_{j=0}^{n-x-2}\{h_{i+j},h_{i+j+x}\}.$ (34)
The detailed proofs and the expression for the Gosset–Mozgunov bound are given in Appendix C.
An important implication of eq. 32 is that the LTI-SDP does not necessarily lower bound the gap of the open boundary system, i.e. $\epsilon_{n}$, in the thermodynamic limit. This is because, the formulation of $\Delta_{\mathrm{LTI}}(n)$ only ensures that it lower bounds the gap of a periodic ring of size $m>2n$. This becomes evident by taking the $n\to\infty$ limit on both sides of eq. 32,
$\lim_{n\to\infty}\Delta_{\mathrm{LTI}}(n)\geq\lim_{n\to\infty}\Delta_{\mathrm{Knabe}}(n)=\lim_{n\to\infty}\epsilon_{n},$ (35)
shows that $\Delta_{\mathrm{LTI}}(n)$ upper bounds $\epsilon_{n}$ in the limit. This is in agreement with the well known fact that gappedness of OBC system implies the gappedness of PBC system, while the converse is not always true. An open question persists as to whether $\Delta_{\mathrm{LTI}}$ converges to the gap in the thermodynamic limit.
IV Numerical benchmarking
In this section we present the results of the numerical application of the method presented in Section II to 1-dimensional spin-chain models. We compared our method with existing finite-size criteria, namely, the Knabe bound (eq. 22) and the bound due to Gosset and Mozgunov (eq. 23).
Because of the inclusions shown in Sec. III, we know that for the same $n$ our method always gives an equal or tighter lower bound on the gap than either of those existing bounds. However, when comparing the methods we should keep in mind that the computational tasks involved are different. While the finite-size methods require the computation of the gap of the $n$-body OBC Hamiltonian—a task with a memory scaling of $\mathcal{O}(d^{n})$—our method optimizes over a positive semidefinite matrix of size $d^{n}\times d^{n}$. The results in this section show that our method outperforms the existing ones even when compared on equal footing (in terms of size of the largest variable in the problem): For the models on which we benchmarked, we observed that $\Delta_{\text{LTI}}(n)\geq\Delta_{\mathrm{Knabe}}(2n)$, and similarly for $\Delta_{\mathrm{GM}}(2n)$.
Another difference between our method and the existing ones is that the finite size criteria assume that the interaction term in the Hamiltonian is a projection. This assumption does not restrict the generality of the methods for the task of detecting a gap, since one can always bound the gap of the original Hamiltonian $\sum_{i}h_{i}$ by the gap of the interaction term $h_{i}$ times the gap of of the projectors Hamiltonian $\sum_{i}\Pi_{i}$, where $\Pi_{i}$ is the projector orthogonal to the ground space of $h_{i}$. However, when the optimization is restricted to be over the new projector Hamiltonian, some information about the exact spectrum of the Hamiltonian is lost in the estimation. Favorably, our method does not need the assumption that local terms have to be projectors. In our numerical benchmarking, we observed situations where better bounds were obtained by solving the LTI SDP with the original Hamiltonian term than with the projector term.
Before presenting the results we address a technical but important issue: When performing the optimization on a computer the SDP solver will produce a solution which is feasible up to some finite precision. If we wish to regard such a solution as providing a mathematical proof of the existence of a gap, we have to make sure that we have a way to obtain a strictly feasible solution from the one produced by the solver. In our case we obtain a numerical solution for the variable $q_{n}$ which should serve as our certificate of a gap (via equation eq. 7). In practice we obtain $q_{n}\succeq-\epsilon$ where $\epsilon$ is the solver precision. For the LTI SDP one cannot simply add $\epsilon\mathbb{I}$ to $q_{n}$ as this would shift the global energy by an extensive amount and a more refined approach has to be taken. In Appendix E, we explain how we modified the LTI SDP to obtain certified solutions.
IV.1 The AKLT model
The AKLT model was the first example of a gapped isotropic spin chain with integral spin [43]. When the model was initially introduced it was proven to be gapped but an explicit lower bound was not provided [44]. In the same year, Knabe introduced his method and provided a lower bound of $0.24806$ for the gap of the AKLT model. Garcia-Saez et al. [45], using a method based on the rotation symmetry of the Hamiltonian, estimated the AKLT gap to be $0.350$, which is compatible with both exact diagonalization results for finite periodic chains (0.35012(0.35012( 0.35012 for $n=16$ [45]), as well as DMRG calculations for finite open chains ($0.35037$ for $n=100$  [^1] Mingru Yang, private communication). Our method proves a lower bound of $\Delta_{\text{AKLT}}\geq 0.34976$ which was obtained by solving the LTI SDP eq. 11) for $n=6$.
Figure 2 shows the Knabe (eq. 22) and Gosset–Mozgunov (eq. 23) lower bounds, and the lower bounds computed with our method (the solution of eq. 11), all for different values of $n$. In addition the finite systems gaps obtained using exact diagonalization for periodic boundary conditions (PBC) are plotted for $n=8,12$ and $16$ as horizontal lines.
We observe that while all three methods (ours, Knabe, and Gosset–Mozgunov) detect the gap starting from system size $n=4$, the LTI SDP performs significantly better than any of the other methods. This is true even if we compare $\Delta_{\text{LTI}}(n)$ with $\Delta_{\mathrm{Knabe}}(2n)$ and $\Delta_{\mathrm{GM}}(2n)$. Recall that the LTI SDP eq. 11 provides a lower bound on the gap of any periodic system of size $\geq 2n$. Our results for the AKLT model show that the LTI SDP gets very close to this optimal value: for $n=6$ we have $\Delta_{\text{PBC}}(12)-\Delta_{\text{LTI}}(6)=0.00036$.
Figure 2: The gap of the $S=1$ AKLT chain. Lower bounds on the energy gap above the ground state of the infinite AKLT chain obtained with different methods are plotted as a function of the size $n$ of the finite system: The Knabe bound eq. 22, the Gosset–Mozgunov bound eq. 23, and our LTI SDP method eq. 11. In addition, gaps of finite systems with periodic boundary conditions (PBC) obtained with exact diagonalization (ED) are plotted for rings of sizes $n=8,12$, and $16$ by dotted, dashed, and solid horizontal black lines, which are almost indistinguishable (the gaps of the $n=12$ and $n=16$ rings differ in the 6th digit). The LTI SDP significantly outperforms the other two methods and gives a lower bound of $0.34976$ for the AKLT gap.
Figure 3: Gaps of the $\mathbb{Z}_{3}$ family of deformed Potts clock models. For a family of spin chain models parameterized by two parameters $r,s$ (eq. 37) lower bounds on the gap were computed at different values of the parameters using different methods: the Knabe and Gosset–Mozgunov finite size criteria (eq. 22 and eq. 23 respectively), and our LTI SDP method eq. 11. The shaded regions indicate the parameter values for which the finite-size methods detected a gap using exact diagonalization results for systems up to size $n=10$ (green region with dashed border for Knabe and blue region with dash-dotted border for Gosset-Mozgunov). The different markers indicate parameter values for which the LTI SDP detected a gap, with different markers corresponding to different values of $n$ (see legend). At each point $r,s$ the marker corresponds to the lowest value of $n$ for which the LTI SDP detected a gap. Darker marker colors indicate larger gaps (see color bar). Each marker corresponds to the larger of the two gaps resulting from using either the original Hamiltonian or the modified Hamiltonian—in which the interaction term is a projection—as the input for the LTI SDP.
IV.2 Deformed clock models
Next we present the results of the application of our method to a family of models generated using a generalization of Witten’s conjugation [30]. In this approach one starts out with a given frustration-free Hamiltonian and deforms it by acting on the local terms with a tensor-product deformation operator. By applying a parameter-dependent deformation one arrives at a family of frustration-free models.
We studied the family of models arising from a deformation of the classical 3-level Potts clock model:
$h_{i,i+1}=2-\sigma_{i}\sigma^{\dagger}_{i+1}-\sigma^{\dagger}_{i}\sigma_{i+1},$ (36)
where $\sigma_{i}$ is an operator that rotates the clock at site $i$ by a third of a turn. Precisely, it is determined by the $\mathbb{Z}_{3}$ clock algebra: $\sigma^{3}=\tau^{3}=\mathbb{I}$, $\sigma^{2}=\sigma^{\dagger}$, $\tau^{2}=\tau^{\dagger}$, and $\sigma\tau=\omega\tau\sigma$, where $\omega=e^{2\pi i/3}$. This model has a gap above its degenerate ground space, spanned by states consisting of tensor products of the eigenstates of $\sigma$ for each of its eigenvalues $\omega^{k},\;k=0,1,2$.
The applied deformation depends on two real parameters $(r,s)$ and the resulting family of models takes the form:
$H(r,s)=$ $-\sum_{i}\sigma_{i}\sigma^{\dagger}_{i+1}+\frac{f(r,s)}{2}(\tau_{j}+\tau_{j+1})+$ (37)
$g_{1}(r,s)\tau_{j}\tau_{j+1}+g_{2}(r,s)\tau_{j}\tau^{\dagger}_{j+1}+\text{h.c.}+\epsilon,$
where the coefficients $f,g_{1}$, and $g_{2}$ are as follows:
$f=-\frac{2}{9}\left[2\left(rs+\omega\frac{r}{s^{2}}+\overline{\omega}\frac{s}{r^{2}}\right)-\left(\frac{1}{rs}+\overline{\omega}\frac{r^{2}}{s}+\omega\frac{s^{2}}{r}\right)\right]$ (38)
$g_{1}=-\frac{2}{9}\left[\omega\left(\frac{r^{2}}{s}+\frac{s}{r^{2}}\right)+\overline{\omega}\left(\frac{s^{2}}{r}+\frac{r}{s^{2}}\right)+rs+\frac{1}{rs}\right]$ (39)
$g_{2}=\frac{1}{9}\left[3+\left(\frac{1}{rs}+\frac{s^{2}}{r}+\frac{r^{2}}{s}\right)-2\left({rs}+\frac{s}{r^{2}}+\frac{r}{s^{2}}\right)\right],$ (40)
and where $\epsilon$ is chosen such that the interaction terms are positive semidefinite.
Due to the fact that the deformation preserves the product-state structure of the ground states of eq. 36, the martingale method can be used to prove that the model is gapped in the whole parameter range [14]: the overlaps between the the ground-state projectors are easy to compute analytically and $\eta$ in eq. 30 is thus shown to decay with the size of the support of the state ($r$ in eq. 30) (see also [47] for a specific treatment of this family of models). This strategy for proving the existence of a gap, however, has the disadvantage that it does not provide us with a concrete lower bound. The exact lower bound proven in this way has a prefactor: the gap of the blocked interaction term $\tilde{h}$ which acts on $2r\gg 1$ spins.
Using Knabe’s criterion, the authors of Ref. [30] computed lower bounds on the gaps of systems with a deformation in a region of the $(r,s)$ parameter space (see Appendix B.2 in  [30]). We benchmarked our method on this family of models to see how it compares to the finite-size criteria in terms of the range of parameters in which a nonzero gap is detected.
We computed lower bounds on the gap for values of $(r,s)$ in the region $(0,1]\times(0,1]$ using the Knabe method (eq. 22), Gosset–Mozgunov (eq. 23) and our method (eq. 11). Figure 3 shows the regions in which each of the methods detected a gap. In the figure, the green-shaded region with dashed boundary indicates where Knabe’s criterion could detect a gap for system sizes $n=3,\ldots,10$, i.e., where $\max_{n\leq 10}\Delta_{\text{Knabe}}(n)>0$; the blue-shaded region with dash-dotted boundary marks the same for the Gosset–Mozgunov bound. The results obtained with our LTI method (eq. 11) are plotted in Figure 3 as individual data points. For each point on a grid in the $(r,s)$ plane we plotted a marker indicating the lowest $n$ for which a positive gap was detected by the LTI SDP (triangle for $n=3$, square for 4, and pentagram for 5).
We observe in Figure 3 that the LTI method for $n=5$ detects a gap in a larger region than Knabe’s bound or the Gosset–Mozgunov bound do for $n=10$ (i.e.  when we compare the methods’ performance for roughly equal computational effort). Furthermore, the region detected by the $n=4$ LTI SDP contains all of the green shaded region (Knabe) and almost all of the blue shaded region (Gosset–Mozgunov).
IV.3 1D Glauber model
Figure 4: Gaps of the 1D Glauber Hamiltonian family. For a family of spin chain models parameterized by two parameters $\gamma,\delta$ (eq. 42) lower bounds on the gap were computed at different values of the parameters using different methods: the Knabe and Gosset–Mozgunov finite size criteria (eq. 22 and eq. 23 respectively) computed for the blocked Hamiltonian $\tilde{h}$ (eq. 46), and our LTI SDP method eq. 11 computed for the original Hamiltonian. The green shaded region (dashed border) marks the region in which the Knabe method detected a gap ($\gamma\lesssim 0.92$), the blue shaded region (dash-dotted border) marks the region where the Gosset–Mozgunov bound did ($\gamma\lesssim 0.96$). The gaps detected by our LTI method are plotted as individual data points with marker colors corresponding to the size of the gap (see color bar).
Glauber models are a family of quantum Hamiltonians related to thermal states of classical Hamiltonians. Under the assumption that these classical thermal states obey the detailed balance condition, they can be mapped to quantum states of the form
|$\ket{\psi}=\frac{1}{\sqrt{Z_{N}}}\sum_{\sigma}e^{-\beta H(\sigma)/2}\ket{\sigma}$ (41)
where $H(\sigma)$ represents a class of classical kinetic Ising models. For this class of Hamiltonians, the state above is found to be representable as matrix product states, thus allowing the construction of a parent Hamiltonian, i.e. a frustration-free Hamiltonian which has the state as its ground state [48]. Here we consider the family of parent Hamiltonians known as the 1D Glauber model [49], which has the local term
$\begin{split}h(\gamma,\delta)=&\mathbb{I}-A(\gamma,\delta)X_{i}+B(\gamma,\delta)Z_{i-1}X_{i}Z_{i+1}\\ &+\delta Z_{i-1}Z_{i+1}-\frac{\gamma(1+\delta)}{2}\left(Z_{i-1}Z_{i}+Z_{i}Z_{i+1}\right),\end{split}$ (42)
where $X,Z$ are Pauli matrices, and $A$, $B$ are scalar coefficients given as
$\begin{split}A(\gamma,\delta)&=\frac{1-\delta}{2}+\frac{(1+\delta)\sqrt{1-\gamma^{2}}}{2},\\ B(\gamma,\delta)&=1-\delta-A(\gamma,\delta).\end{split}$ (43)
The model approaches criticality for $\gamma\to 1$, and the scaling behavior of the gap in this regime allows for the study of dynamical exponents [49]. Additionally, it can be verified that the eigenvectors of $h(\gamma,\delta)$ are only dependent on $\gamma$, and the eigenvalues are
$0,0,0,0,2(1-\delta),2(1-\delta),2(1+\delta),2(1+\delta),$ (44)
i.e., independent of $\gamma$. Note that for calculating finite size criteria bounds, the local gap threshold $\epsilon_{n}$ (Sec. III) corresponds to the gap of the projector Hamiltonian, which is independent of $\delta$.
Finally, to compute the finite size criteria bounds for a 3-local Hamiltonian, we have to adapt the calculation of $\Delta_{\mathrm{Knabe}}$ and $\Delta_{\mathrm{GM}}$ using a two-step blocking procedure. We first block two $\mathrm{spin}-\frac{1}{2}$ particles of the original chain into one $\mathrm{spin}-\frac{3}{2}$ $(d=4)$ particle,
$\{i,i+1\}$$\rightarrow\mathcal{I}.$ (45)
In the next step, we combine consecutive local terms of the original Hamiltonian (eq. 42) into one term,
$\tilde{h}_{\mathcal{I},\mathcal{I}+1}$ $=h_{i,i+1}+h_{i+1,i+2},$ (46)
which provides the expression for the nearest neighbor Hamiltonian $\tilde{h}$ on the new $\mathrm{spin}-\frac{3}{2}$ chain.
Figure 4 shows the parameter region in which the finite size criteria detect a gap compared to the region in which the LTI SDP method does. For the range $\gamma,\delta\in[0,1)$, we computed the Knabe and Gosset–Mozgunov bounds for a system of 6 blocked spins ($n=12$ of the original spins) with interaction term $\tilde{h}$; as well as the LTI SDP for $n=6$ with the original Hamiltonian from eq. 42. For the LTI results the color of the marker at each point corresponds to the size of the detected gap (see color bar).
We observe that the LTI SDP allows us to detect a gap orders of magnitude closer to the critical value ($\gamma=1$) than what is possible using the two other finite size criteria.
V Conclusion and Outlook
In this paper, we put forward a semidefinite programming hierarchy for certifying the spectral gap of frustration-free Hamiltonians. The method relies on a well-known relation between the positivity of a quadratic operator in the Hamiltonian and the existence of a spectral gap. Each level $n$ in our hierarchy searches for the maximal value $\Delta_{\mathrm{LTI}}$ for which the quadratic operator on the global system can be decomposed as a sum of $n$-local positive terms. Thus, it involves an $n$-body optimization problem that can be efficiently solved with semidefinite programming.
We proved that our method matches or outperforms the most widely used finite-size criteria like the Knabe [15] or Gosset–Mozgunov [20] bounds. Extensive numerical benchmarks on 1D frustration-free models show that our method significantly outperforms the above criteria, when compared at the same computational cost. In particular, for the spectral gap of the AKLT chain, we provide bounds that tightly match the best known estimates. For other families of models we demonstrated that our method detects gaps in a significantly larger parameter range than previous finite size criteria do, especially in models close to criticality.
Interestingly, the dual formulation of our problem corresponds to the minimization of the first excited energy on marginals satisfying the so-called local translational invariance (LTI) property. Those marginals can be seen as a relaxation of the first excited reduced states. In this sense, our method adds to the recent advancements on SDP relaxations to estimate the ground state [40, 50], and to compute its physical properties [51].
An interesting question we did not address in this paper is the completeness of the LTI hierarchy. Proving that the lower bound produced by our method converges to the infinite-system gap as $n$ is increased is left for future work. It would also be interesting to obtain a bound on the convergence rate to the exact spectral gap in the thermodynamic limit.
We believe it is possible to adapt our method to the case of systems with open boundary conditions. It is known that finite-size criteria can be modified to take into account edge terms in the Hamiltonian [22], such as those in the Motzkin and Fredkin chains [52, 53]. We expect that, in a similar fashion, one can modify the LTI SDP to allow for the treatment of such systems as well.
Finally, from its formulation, we believe that our method can allow the certification of gaps for systems in higher spatial dimensions. The generalization of our method to such systems is straightforward. The challenge in applying our method in this setting lies solely in the ability to solve large scale SDPs. Yet, from the results we presented for one-dimensional systems it seems reasonable to expect that the system size for which one would need to solve the problem should be significantly smaller than the system sizes involved in proving gaps using finite size criteria in 2D (e.g. [34]).
Acknowledgements.
We thank Milán Rozmán for guiding us through the martingale literature. We thank Mingru Yang for enlightening discussions as well as for providing us with the DMRG result for the AKLT model. We thank Ignacio Cirac for stimulating discussions at the beginning of the project. We thank Jin-Fu Chen for useful discussions about the Glauber model. I.K. and N.S. acknowledge support from the European Union’s Horizon 2020 research and innovation programme through Grant No. 863476 (ERC-CoG SEQUAM), and the FWF Quantum Austria Funding Initiative project “Entanglement Order Parameters” (Grant DOI 10.55776/P36305), funded through the European Union – NextGenerationEU. N.S. also acknowledges support by the the Austrian Science Fund FWF through Grant DOIs 10.55776/COE1 and 10.55776/F71. K.S.R. and J.T. acknowledge the support received from the European Union’s Horizon Europe research and innovation programme through the ERC StG FINE-TEA-SQUAD (Grant No. 101040729). P.E. and J.T. acknowledge the support received by the Dutch National Growth Fund (NGF), as part of the Quantum Delta NL programme. P.E. acknowledges the support received through the NWO-Quantum Technology programme (Grant No. NGF.1623.23.006). This publication is part of the ‘Quantum Inspire - the Dutch Quantum Computer in the Cloud’ project (with project number [NWA.1292.19.194]) of the NWA research program ‘Research on Routes by Consortia (ORC)’, which is funded by the Netherlands Organization for Scientific Research (NWO). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the funding institutions. Neither of the funding institutions can be held responsible for them. Parts of this work were performed by using the compute resources from the Academic Leiden Interdisciplinary Cluster Environment (ALICE) provided by Leiden University. F.B. acknowledges financial support from the Alexander von Humboldt Foundation.
Appendix A The LTI condition is the only freedom in TI Hamiltonians
In the main text, we used the property that if two local operators supported on $n$ sites, satisfy the following condition
$\sum_{i}q_{n,i}=\sum_{i}g_{n,i}$ (47)
then they are related as (eq. 10 of main text)
$q_{n,i}=g_{n,i}+\mathbb{I}\otimes Y-Y\otimes\mathbb{I}.$ (48)
Now, we will formally derive this property. Let $X$ be an $n$-body operator. Claim:
$\sum_{i}\tau_{i}(X)=0\iff\exists A\text{ such that }X=A\otimes\mathbb{I}-\mathbb{I}\otimes A.$ (49)
Here $\tau_{i}(X)$ denotes the translation of operator $X$ by $i$ sites.
Proof: For the direction $\mathrm{RHS}\Rightarrow\mathrm{LHS}$, notice that the terms $A\otimes\mathbb{I}-\mathbb{I}\otimes A$ cancel telescopically, meaning that
$\tau_{i}(A\otimes\mathbb{I}-\mathbb{I}\otimes A)=0.$ (50)
It remains to prove the other implication, i.e.,
$\tau_{i}(X)\Rightarrow X=A\otimes\mathbb{I}-\mathbb{I}\otimes A.$ (51)
To prove this, we take an orthonormal basis consisting of strings of single site operators $S_{i_{1},i_{2},\ldots,i_{n}}:=\sigma_{i_{1}}\otimes\sigma_{i_{2}}\otimes\ldots\sigma_{i_{n}}$ with $\sigma_{0}$ proportional to the identity $\mathbb{I}$. Expanding $X$ in this basis,
$X=\sum_{i_{1},i_{2},\ldots,i_{n}}X_{i_{1},i_{2},\ldots,i_{n}}S_{i_{1},i_{2},\ldots,i_{n}}.$ (52)
In the above, first consider the basis elements with $i_{1}\neq 0$ and $i_{n}\neq 0$ and observe that in
$\textrm{Tr}\left(\sum_{j\in\mathbb{Z}}\tau_{j}(X)S_{i_{1},i_{2},\ldots,i_{n}}\right),$ (53)
we get that only one term in the sum survives, i.e. the term with $j=1$. Thus, $X$ satisfies
$X_{i_{1},i_{2},\ldots,i_{n}}=0\text{ for }i_{1},i_{n}\neq 0\text{ and for all }i_{2},\ldots,i_{n-1}=0,\ldots,d^{2}-1$ (54)
In the next step, we observe that $S_{0,i_{2},\ldots,i_{n}}$ with $i_{2}\neq 0$ and $i_{n}\neq 0$, which implies
$X_{0,i_{2},\ldots,i_{n}}+X_{i_{2},\ldots,i_{n},0}=0,$ (55)
i.e.
$X_{0,i_{2},\ldots,i_{n}}=-X_{i_{2},\ldots,i_{n},0},$ (56)
showing that $X$ has components $A^{(2)}\otimes\mathbb{I}-\mathbb{I}\otimes A^{(2)}$ with $A^{(2)}=\sum X_{i_{2},\ldots,i_{n},0}S_{i_{2},\ldots,i_{n}}$.
Continuing as before, we now observe that for $S_{0,0,i_{3},\ldots,i_{n}}$ with $i_{3}\neq 0$ and $i_{n}\neq 0$, we get
$X_{0,0,i_{3},\ldots,i_{n}}+X_{0,i_{3},\ldots,i_{n},0}+X_{i_{3},\ldots,i_{n},0,0}=0.$ (57)
Defining the $(n-2)$-body operators $B:=\sum_{i_{3},\ldots,i_{n}}X_{i_{3},\ldots,i_{n},0,0}S_{i_{3},\ldots,i_{n}}$, and $C:=\sum_{i_{3},\ldots,i_{n}}X_{0,i_{3},\ldots,i_{n},0}S_{i_{3},\ldots,i_{n}}$, we note that $X$ has contributions
$X=\ldots+$ $\sum_{i_{3},\ldots,i_{n}}X_{i_{3},\ldots,i_{n},0,0}S_{i_{3},\ldots,i_{n}}\otimes\mathbb{I}\otimes\mathbb{I}+$ (58)
$\sum_{i_{3},\ldots,i_{n}}X_{0,i_{3},\ldots,i_{n},0}\mathbb{I}\otimes S_{i_{3},\ldots,i_{n}}\otimes\mathbb{I}+$
$\sum_{i_{3},\ldots,i_{n}}(-X_{i_{3},\ldots,i_{n},0,0}-X_{0,i_{3},\ldots,i_{n},0})\mathbb{I}\otimes\mathbb{I}\otimes S_{i_{3},\ldots,i_{n}}=$
$=$ $\ldots+B\otimes\mathbb{I}\otimes\mathbb{I}+\mathbb{I}\otimes C\otimes\mathbb{I}-\mathbb{I}\otimes\mathbb{I}\otimes(B+C)$
Now defining $B^{\prime}=B$, $C^{\prime}=B+C$, we can rewrite the last line in eq. 58 as
$X=\ldots+B^{\prime}\otimes\mathbb{I}\otimes\mathbb{I}+\mathbb{I}\otimes(C^{\prime}-B^{\prime})\otimes\mathbb{I}-\mathbb{I}\otimes\mathbb{I}\otimes(C^{\prime})=\ldots+A^{(3)}\otimes\mathbb{I}-\mathbb{I}\otimes A^{(3)}$ (59)
with $A^{(3)}:=B^{\prime}\otimes\mathbb{I}+\mathbb{I}\otimes C^{\prime}$.
Repeating the same procedure for all $k\leq n$: Let $i_{k},i_{n}\neq 0$ and take the trace of $\sum_{j}\tau_{j}(X)$ with $S_{0,0,\ldots,0,i_{k},\ldots,i_{n}}$. Similarly as in the eq. 57 we obtain a condition which involves $k$ types of coefficients of $X$:
$X_{i_{k},\ldots,i_{n},0,\ldots,0}+X_{0,i_{k},\ldots,i_{n},\ldots,0}+X_{0,\ldots,i_{k},\ldots,i_{n},0}+\ldots+X_{0,\ldots,i_{k},\ldots,i_{n}}=0,$ (60)
which lead to contributions to $X$ of the form:
$X=\ldots+$ $T_{1}\otimes\mathbb{I}\otimes\mathbb{I}\otimes\ldots\mathbb{I}+$ (61)
$\mathbb{I}\otimes T_{2}\otimes\mathbb{I}\otimes\ldots\mathbb{I}+$
$\vdots$
$\mathbb{I}\otimes\mathbb{I}\otimes\mathbb{I}\otimes\ldots T_{k},$
where each $T_{i}$ is an $(n-k+1)$-body operator. Now for $l=1,\ldots,k-1$, we define the terms $T_{l}^{\prime}=\sum_{i=1}^{l}T_{i}$ and use $T_{k}=-\sum_{i=1}^{k-1}T_{i}=-T^{\prime}_{k-1}$ from eq. 60 to rewrite the terms in eq. 61 as $A^{(k)}\otimes\mathbb{I}-\mathbb{I}\otimes A^{(k)}$ where
$A^{(k)}:=[T_{1}^{\prime}\otimes\mathbb{I}\otimes\ldots\mathbb{I}+\mathbb{I}\otimes T_{2}^{\prime}\otimes\ldots\mathbb{I}+\ldots\mathbb{I}\otimes\mathbb{I}\otimes\ldots T_{k-1}^{\prime}].$ (62)
Finally collecting all the coefficients of $X$, we obtain that $X=A\otimes\mathbb{I}-\mathbb{I}\otimes A$, where $A=\sum_{k=1}^{n}A^{(k)}$.
Appendix B First excited state is the optimal solution $\rho$ of the dual program
We take the dual objective function of the exact dual SDP eq. 15, and expand it in the eigenbasis of $H$,
[…]
$=\sum_{j}E_{j}^{2}\rho_{jj}$ (64)
$=\Delta^{2}\rho_{11}+\sum\limits_{j\geq 2}E_{j}^{2}\rho_{jj}.$ (65)
Similarly, expanding the constraint,
[…]
$=\Delta\rho_{11}+\sum\limits_{j\geq 2}E_{j}\rho_{jj}=1$ (67)
[…]
$=\Delta+\sum\limits_{j\geq 2}(E_{j}^{2}-\Delta E_{j})\rho_{jj},$ (69)
Since $E_{j}^{2}-\Delta E_{j}\geq 0\quad\forall\,j\geq 2$, the minimum is attained for $\rho_{jj}=0\quad\forall\,j\geq 2$. At the optimal $\rho^{*}$,
[…]
Now using positivity constraint on $\rho$, we find that
$\rho_{ij}=0\quad\forall\,\,i\neq j.$ (71)
which shows that the dual variable converges to the reduced marginals of the first excited state.
Appendix C Proof for inclusion of finite-size based methods
We show that the finite size criteria bounds arise as a special case of the SDP eq. 11. The operator $Q_{n}(\delta)$ which is a truncation of $H^{2}-\delta H$ is defined as,
$\begin{split}Q_{n}(\delta_{n})&=\sum_{i=1}^{N}h_{i}^{2}+\sum_{i}\left(\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\}\right)-\delta_{n}\sum_{i=1}^{N}h_{i},\end{split}$ (72)
where $\delta_{n}$ can take any value in the feasible set of the SDP in eq. 11. For this we need to find a positive generating local term $q_{n}(\delta)\succeq 0$ such that,
$Q_{n}(\delta_{n})$ $=\sum_{i}q_{n,i}(\delta)\quad\text{ and }\quad q_{n,i}(\delta)\succeq 0.$ (73)
In the following subsections, we derive $q_{n}(\delta)$ for the bound $\delta$ given by (1) Knabe bound [15], (2) Gosset–Mozgunov bound [20], (3) Martingale bound [14]. Following the assumptions of these methods, in the proof, we focus on the case of nearest neighbor Hamiltonians, and assume that the local terms are projectors, i.e. $h_{i}^{2}=h_{i}$.
C.1 Knabe bound eq. 22
The finite size criteria bounds are based on comparing the contribution of different overlapping and non-overlapping anticommutator terms $\{h_{i},h_{j}\}$ in $H^{2}$. The method is based on deriving an operator approximation for global operator $H^{2}$ (on the full system), in terms of the open boundary Hamiltonian on a finite system of $n$ sites. We define the open boundary Hamiltonian as
$h_{n,i}=\sum_{j=i}^{i+n-2}h_{j,j+1}.$ (74)
where the index $i$ denotes the index of the starting site. Note that this Hamiltonian is supported on sites $i$ to $i+n-1$. The open boundary Hamiltonian satisfies the following relation,
$h_{n,i}^{2}\succeq\epsilon_{n}h_{n,i},$ (75)
where $\epsilon_{n}$ is a lower bound on the gap of $h_{n,i}$. Then, the Knabe bound is given as
$\Delta_{m}^{\circ}\geq\Delta_{\mathrm{Knabe}}(n)=\frac{n-1}{n-2}\left(\epsilon_{n}-\frac{1}{n-1}\right).$ (76)
It provides a direct relation between the finite size gap $\epsilon_{n}$ and the gap of system of size $m>n$. To prove the inclusion of the Knabe bound, we require a positive operator decomposition for $Q_{n}(\Delta_{\mathrm{Knabe}}(n))$. The first step is to write the expansion of $h_{n,i}^{2}$,
$h_{n,i}^{2}$ $=\left(h_{i}+...+h_{i+n-2}\right)^{2}$ (77)
$=h_{i}^{2}+...+h_{i+n-2}^{2}+\sum_{j=0}^{n-3}\{h_{i+j},h_{i+j+1}\}+\sum_{j=0}^{n-4}\{h_{i+j},h_{i+j+2}\}+...+\{h_{i},h_{i+n-2}\}$ (78)
$=h_{n,i}+\sum_{j=0}^{n-3}\{h_{i+j},h_{i+j+1}\}+\sum_{j=0}^{n-4}\{h_{i+j},h_{i+j+2}\}+...+\{h_{i},h_{i+n-2}\}$ (79)
where in the last step, we used that the local terms are projectors $h_{i}^{2}=h_{i}$. Summing on the left and right over all sites $i$,
$\sum_{i=1}^{N}h_{n,i}^{2}$ $=(n-1)H+(n-2)\sum_{i}\{h_{i},h_{i+1}\}+(n-3)\sum_{i}\{h_{i},h_{i+2}\}+...+\sum_{i}\{h_{i},h_{i+n-2}\}.$ (80)
Now we will derive a positive operator decomposition for $Q_{n}(\Delta_{\mathrm{Knabe}}(n))$. First let us write the operator
$Q_{n}(\Delta_{\mathrm{Knabe}}(n))$ $=H+\sum_{i}\left(\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\}\right)-\Delta_{\mathrm{Knabe}}(n)H,$ (81)
Substituting the Knabe bound eq. 76 into the above equation,
$Q_{n}(\Delta_{\mathrm{Knabe}}(n))$ $=H+\sum_{i}\left(\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\}\right)-\left(\frac{n-1}{n-2}\epsilon_{n}-\frac{1}{n-2}\right)H,$ (82)
$=\left(\frac{1-\epsilon_{n}}{n-2}\right)(n-1)H+\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\}$ (83)
Now, we notice that $(n-1)H$ can be expanded as
$(n-1)H=\sum_{i}h_{i}+h_{i+1}+...+h_{i+n-2},$ (84)
Substituting this in eq. 83,
$Q_{n}(\Delta_{\mathrm{Knabe}}(n))$ $=\sum_{i}\left(\frac{1-\epsilon_{n}}{n-2}\right)(h_{i}+h_{i+1}+...+h_{i+n-2})+\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\}.$ (85)
In the above expression, $Q_{n}$ has a $n-$local decomposition. However, we require a positive $n-$local operator decomposition $q_{n}$, which can be obtained by rearranging the terms as allowed by the LTI freedom (eq. 12, App. A). For the anticommutators, we use the following replacements,
$\{h_{i},h_{i+1}\}$ $\rightarrow\frac{1}{n-2}\sum_{j=0}^{n-3}\{h_{i+j},h_{i+j+1}\}$ (86)
$\{h_{i},h_{i+2}\}$ $\rightarrow\frac{1}{n-3}\sum_{j=0}^{n-4}\{h_{i+j},h_{i+j+2}\}$ (87)
$\vdots$
$\{h_{i},h_{i+n-3}\}$ $\rightarrow\frac{1}{2}\left(\{h_{i},h_{i+2}\}+\{h_{i+1},h_{i+3}\}\right)$ (88)
which when summed over $i$ on both sides, recover the same global operator. For example, in the first replacement, the two sides are locally related by an operator $Y$ supported on $n-1$ sites, which satisfies the equation,
$\{h_{i},h_{i+1}\}+\mathbb{I}\otimes Y-Y\otimes\mathbb{I}=\frac{1}{n-2}\sum_{j=0}^{n-3}\{h_{i+j},h_{i+j+1}\}.$ (89)
Using the above replacements, the new form for $Q_{n}$ is,
$Q_{n}(\Delta_{\mathrm{Knabe}}(n))$ $=\sum_{i}\left[\left(\frac{1-\epsilon_{n}}{n-2}\right)h_{n,i}+\frac{1}{n-2}\sum_{j=0}^{n-3}\{h_{i+j},h_{i+j+1}\}+\frac{1}{n-3}\sum_{j=0}^{n-4}\{h_{i+j},h_{i+j+2}\}+...+\{h_{i},h_{i+n-2}\}\right]$ (90)
$:=\sum_{i}q_{n,i}(\Delta_{\mathrm{Knabe}}(n)).$ (91)
where above we chose a $n-$local form for $q_{n}$. We need to show that
$q_{n,i}(\Delta_{\mathrm{Knabe}}(n)):=\left(\frac{1-\epsilon_{n}}{n-2}\right)h_{n,i}+\frac{1}{n-2}\sum_{j=0}^{n-3}\{h_{i+j},h_{i+j+1}\}+\frac{1}{n-3}\sum_{j=0}^{n-4}\{h_{i+j},h_{i+j+2}\}+...+\{h_{i},h_{i+n-2}\}$ (92)
is positive semidefinite, i.e. $q_{n,i}(\Delta_{\mathrm{Knabe}}(n))\overset{!}{\succeq}0$. Using $h_{n,i}^{2}\succeq\epsilon_{n}h_{n,i}$,
$q_{n,i}(\Delta_{\mathrm{Knabe}}(n))$ $\succeq\frac{1}{n-2}(h_{n,i}-h_{n,i}^{2})+\frac{1}{n-2}\sum_{j=0}^{n-3}\{h_{i+j},h_{i+j+1}\}+\frac{1}{n-3}\sum_{j=0}^{n-4}\{h_{i+j},h_{i+j+2}\}+...+\{h_{i},h_{i+n-2}\}$ (93)
Substituting $h_{n,i}^{2}$ from eq. 79, we obtain
$\begin{split}q_{n,i}(\Delta_{\mathrm{Knabe}}(n))&\succeq\frac{1}{(n-3)(n-2)}\sum_{j=0}^{n-4}\{h_{i+j},h_{i+j+2}\}+\frac{2}{(n-4)(n-2)}\sum_{j=0}^{n-5}\{h_{i+j},h_{i+j+3}\}...+\frac{n-3}{n-2}\{h_{i},h_{i+n-2}\}\\ &\succeq 0,\end{split}$ (94)
where the positivity follows because the RHS is sum of disjoint anticommutators. Thus, we explicitly found a $n-$local positive operator decomposition for $[H^{2}]_{n}-\Delta_{\mathrm{Knabe}}(n)$, proving that $\Delta_{\mathrm{Knabe}}(n)$ lies in the feasible set of the LTI SDP.
C.2 Gosset–Mozgunov bound eq. 23
Similar to the previous section, we will derive the positive $n-$local term $q_{n}(\Delta_{\mathrm{GM}}(n))$ for the Gosset–Mozgunov bound which has the expression
$\Delta_{m}^{\circ}\geq\Delta_{\mathrm{GM}}(n)=\frac{5}{6}\frac{n^{2}+n}{n^{2}-4}\left(\epsilon_{n}-\frac{6}{n^{2}+n}\right),$ (95)
where it lower bounds the gap of system size $m>2n$. Expanding the global operator $Q_{n}(\Delta_{\mathrm{GM}}(n))$, we obtain
$Q_{n}(\Delta_{\mathrm{GM}}(n))$ $=H+\{h_{i},h_{i+1}\}+...+\{h_{i},h_{i+n-2}\}-\Delta_{\mathrm{GM}}(n)H,$ (96)
$=(1-\Delta_{\mathrm{GM}}(n))H+\sum_{i}\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\}$ (97)
Now, let us simplify the coefficient $(1-\Delta_{\mathrm{GM}}(n))$ of the first term as,
$1-\Delta_{\mathrm{GM}}(n)=a-b\epsilon_{n},$ (98)
where we defined
$a=\frac{n^{2}+1}{n^{2}-4},\quad b=\frac{5}{6}\frac{n^{2}+n}{n^{2}-4}$ (99)
for better readability of the upcoming steps. In eq. 97,
$Q_{n}(\Delta_{\mathrm{GM}}(n))$ $=\sum_{i}(ah_{i}-b\epsilon_{n}h_{i})+\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\}$ (100)
The following discussion heavily uses notation from Ref. [20]. Firstly, define
$B_{n,i}=\sum_{j=0}^{n-2}c_{j}h_{i+j}$ (101)
where coefficients $c_{j}$ satisfy the relations from equations (5)-(7) of Ref. [20]. Then, expanding $B_{n,i}^{2}$, we obtain
$B_{n,i}^{2}$ $=\sum_{j=0}^{n-2}c_{j}^{2}h_{i+j}+\sum_{j=0}^{n-3}c_{j}c_{j+1}\{h_{i+j},h_{i+j+1}\}+\sum\limits_{j,k\in\{0,..,n-2\};|j-k|\geq 2}c_{j}c_{k}h_{i+j}h_{i+k}.$ (102)
Using the LTI freedom (App. A), we can perform the following replacement of $h_{i}$ in the first term eq. 100 with
$h_{i}\to\frac{1}{\sum_{r=0}^{n-2}c_{r}^{2}}\sum_{j=0}^{n-2}c_{j}^{2}h_{i+j}$ (103)
and the $h_{i}$ in the second term of eq. 100, with
$h_{i}$ $\to\frac{B_{n,i}}{\sum_{r=0}^{n-2}c_{r}}.$ (104)
Note that as for the previous proof for the Knabe bound, we can perform these replacements (using LTI condition) because they do not modify the complete sum over the index $i$. Substituting the above two equations into eq. 100,
$Q_{n}(\Delta_{\mathrm{GM}}(n))$ $=\sum_{i}\left(\sum_{j=0}^{n-2}\frac{ac_{j}^{2}}{\sum_{r=0}^{n-2}c_{r}^{2}}h_{i+j}-b\epsilon_{n}\frac{B_{n,i}}{\sum_{r=0}^{n-2}c_{r}}\right)+\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\}.$ (105)
The next step is to use the LTI freedom to change the local form of the anticommutators,
$\{h_{i},h_{i+1}\}$ $\to\frac{\sum_{j=0}^{n-3}c_{j}c_{j+1}\{h_{i+j},h_{i+j+1}\}}{\sum_{r=0}^{n-3}c_{r}c_{r+1}}$ (106)
$\sum_{x=2}^{n-2}\{h_{i},h_{i+x}\}$ $\to\frac{\sum\limits_{j,k\in\{0,..,n-2\};|j-k|\geq 2}c_{j}c_{k}h_{i+j}h_{i+k}}{\sum\limits_{q,r\in\{0,..,n-2\};|q-r|\geq 2}c_{q}c_{r}}.$ (107)
keeping the global sum of eq. 100 the same. Substituting the above two equations into eq. 105,
$\begin{split}Q_{n}(\Delta_{\mathrm{GM}}(n))=\sum_{i}\left(\sum_{j=0}^{n-2}\frac{ac_{j}^{2}}{\sum_{r=0}^{n-2}c_{r}^{2}}h_{i+j}-b\epsilon_{n}\frac{B_{n,i}}{\sum_{r=0}^{n-2}c_{r}}\right)&+\frac{\sum_{j=0}^{n-3}c_{j}c_{j+1}\{h_{i+j},h_{i+j+1}\}}{\sum_{r=0}^{n-3}c_{r}c_{r+1}}\\ &+\frac{\sum\limits_{j,k\in\{0,..,n-2\};|j-k|\geq 2}c_{j}c_{k}h_{i+j}h_{i+k}}{\sum\limits_{q,r\in\{0,..,n-2\};|q-r|\geq 2}c_{q}c_{r}}.\end{split}$ (108)
Now, using the above expansion, we define $q_{n}(\Delta_{\mathrm{GM}}(n))$ as,
$\begin{split}q_{n}(\Delta_{\mathrm{GM}}(n)):=\left(\sum_{j=0}^{n-2}\frac{ac_{j}^{2}}{\sum_{r=0}^{n-2}c_{r}^{2}}h_{i+j}-b\epsilon_{n}\frac{B_{n,i}}{\sum_{r=0}^{n-2}c_{r}}\right)&+\frac{\sum_{j=0}^{n-3}c_{j}c_{j+1}\{h_{i+j},h_{i+j+1}\}}{\sum_{r=0}^{n-3}c_{r}c_{r+1}}\\ &+\frac{\sum\limits_{j,k\in\{0,..,n-2\};|j-k|\geq 2}c_{j}c_{k}h_{i+j}h_{i+k}}{\sum\limits_{q,r\in\{0,..,n-2\};|q-r|\geq 2}c_{q}c_{r}}.\end{split}$ (109)
And the requirement is to prove the positivity for the above choice of
$q_{n}(\Delta_{\mathrm{GM}}(n))\overset{!}{\succeq}0.$ (110)
The first step is to use the result of Lemma 4 of Ref. [20], which is the relation
$B_{n,i}^{2}\succeq\epsilon_{n}\left(\frac{1}{n-1}\sum_{j=0}^{n-2}c_{j}\right)B_{n,i}.$ (111)
Secondly, we replace the $B_{n,i}$ in the second term of eq. 109 with $B_{n,i}^{2}$ using the above relation,
$q_{n}(\Delta_{\mathrm{GM}}(n))$ $\succeq\frac{a\sum_{j=0}^{n-2}c_{j}^{2}h_{i+j}}{\sum_{r=0}^{n-2}c_{r}^{2}}-\frac{b(n-1)B_{n,i}^{2}}{\left(\sum_{r=0}^{n-2}c_{r}\right)^{2}}+\frac{\sum_{j=0}^{n-3}c_{j}c_{j+1}\{h_{i+j},h_{i+j+1}\}}{\sum_{r=0}^{n-3}c_{r}c_{r+1}}+\frac{\sum\limits_{j,k\in\{0,..,n-2\};|j-k|\geq 2}c_{j}c_{k}h_{i+j}h_{i+k}}{\sum\limits_{q,r\in\{0,..,n-2\};|q-r|\geq 2}c_{q}c_{r}}$ (112)
$=\sum_{j=0}^{n-2}\left(\frac{a}{\sum_{r=0}^{n-2}c_{r}^{2}}-\frac{b(n-1)}{\left(\sum_{r=0}^{n-2}c_{r}\right)^{2}}\right)c_{j}^{2}h_{i+j}+\left(\frac{1}{\sum_{r=0}^{n-3}c_{r}c_{r+1}}-\frac{b(n-1)}{\left(\sum_{r=0}^{n-2}c_{r}\right)^{2}}\right)\sum_{j=0}^{n-3}c_{j}c_{j+1}\{h_{i+j},h_{i+j+1}\}$ (113)
$\quad\quad\quad\quad\quad\quad\quad+\sum\limits_{j,k\in\{0,..,n-2\};|j-k|\geq 2}\left(\frac{1}{\sum\limits_{q,r\in\{0,..,n-2\};|q-r|\geq 2}c_{q}c_{r}}-\frac{b(n-1)}{\left(\sum_{r=0}^{n-2}c_{r}\right)^{2}}\right)c_{j}c_{k}h_{i+j}h_{i+k}$
There are three terms based on $c_{j}$’s, that can be simplified using relations from equations (30)-(32) of Ref. [20], which are
$\sum_{r=0}^{n-2}c_{r}$ $=\frac{n^{3}-n}{6}$ (114)
$\sum_{r=0}^{n-2}c_{r}^{2}$ $=\frac{n^{5}-n}{30}$ (115)
$\sum_{r=0}^{n-3}c_{r}c_{r+1}$ $=\frac{n^{5}-5n^{3}+4n}{30}.$ (116)
Substituting the above relations into eq. 113, we obtain that the coefficient of second term is exactly zero, i.e.
$\left(\frac{1}{\sum_{j=0}^{n-3}c_{j}c_{j+1}}-\frac{b(n-1)}{\left(\sum_{j=0}^{n-2}c_{j}\right)^{2}}\right)=0.$ (117)
Similarly, upon simplifying the coefficient of the first term,
$\left(\frac{a}{\sum_{r=0}^{n-2}c_{r}^{2}}-\frac{b(n-1)}{\left(\sum_{r=0}^{n-2}c_{r}\right)^{2}}\right)=30\left(\frac{(n^{2}+1)}{n(n^{4}-1)(n^{2}-4)}-\frac{n^{2}+1}{n(n^{4}-1)(n^{2}-4)}\right)=0,$ (118)
which is also zero. Finally, the coefficient of the last term is simplified using the autocorrelation lemma from Ref. [20],
$\sum\limits_{j,k\in\{0,..,n-2\};|j-k|\geq 2}c_{j}c_{k}\leq\sum_{j=0}^{n-3}c_{j}c_{j+1}.$ (119)
Substituting it in the last term of eq. 113, we find
$\left(\frac{1}{\sum\limits_{j,k\in\{0,..,n-2\};|j-k|\geq 2}c_{j}c_{k}}-\frac{b(n-1)}{\left(\sum_{j=0}^{n-2}c_{j}\right)^{2}}\right)\geq\left(\frac{1}{\sum_{j=0}^{n-3}c_{j}c_{j+1}}-\frac{b(n-1)}{\left(\sum_{j=0}^{n-2}c_{j}\right)^{2}}\right)=0.$ (120)
Combining all three terms, we finally obtain
$q_{n}(\Delta_{\mathrm{GM}}(n))\succeq 0.$ (121)
Thus, we explicitly derived a $n-$local positive operator decomposition for $[H^{2}]_{n}-\Delta_{\mathrm{GM}}(n)$, proving that $\Delta_{\mathrm{GM}}(n)$ lies in the feasible set of the LTI SDP.
C.3 Martingale condition bound eq. 31
We show that the martingale bound of the blocked Hamiltonian $K$ is contained in the feasible set of the LTI SDP. We will mainly use the inequality (Lemma 6.3(2) of [16])
$\{P_{\mathcal{I},\mathcal{I}+1},P_{\mathcal{I}+1,\mathcal{I}+2}\}\succeq-\eta(P_{\mathcal{I},\mathcal{I}+1}+P_{\mathcal{I}+1,\mathcal{I}+2}),$ (122)
The lower bound on the gap of Hamiltonian $K$ after using the above inequality is
$K^{2}\succeq(1-2\eta)K.$ (123)
We take that $K$ is supported on $n=3r$ sites of the original Hamiltonian. Therefore, to prove that the above bound is included in the feasible set, we find a decomposition for $Q_{n}(1-2\eta)$ such that
$q_{n}(1-2\eta)\overset{!}{\succeq}0.$ (124)
We start by expanding the LHS,
$Q_{n}(1-2\eta)$ $=H+\sum_{i}(\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\})-(1-2\eta)H$ (125)
$=H+\sum_{i}(\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\})-(1-2\eta)H$ (126)
$=\sum_{i}2\eta h_{i}+(\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\})$ (127)
$=\sum_{i}\eta(h_{i}+h_{i+1})+(\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\})$ (128)
Defining the local term to be
$q_{n,i}(1-2\eta):=\eta(h_{i}+h_{i+1})+(\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\})$ (129)
Using the relation eq. 122, we obtain that
$q_{n,i}(1-2\eta)\succeq 0.$ (130)
Appendix D Optimizing the interaction term using the SDP gradient formula
It is known that the gradient of the optimal objective of an SDP with respect to the problem data can be obtained from the primal-dual pair of variables attaining the optimal solution [39].
First of all, we see that upon applying the SDP gradient formula ([39], Section 4.2) to the exact SDP formulation of the gap problem section II, we recover the result of the Hellmann–Feynman theorem: Let $H$ be the full system Hamiltonian with ground-state energy zero, and let $P_{0}$ be the projector to the global ground state. Assume $H$ depends on a parameter $x$ as $H(x)=H_{0}+(\mathbb{I}-P_{0})A(x)(\mathbb{I}-P_{0})$, where $A$ is some operator depending smoothly on $x$, such that $H(x)$ has the same ground state as $H_{0}$ in a neighborhood of $x_{0}$.
The SDP gradient formula tells us that the derivative of the objective function (in our case the exact gap) with respect to the parameter $x$ is given by the partial derivative of the Lagrangian at the primal-dual optimal point. In the same spirit as the Hellmann–Feynman theorem—only the explicit dependence on $x$ contributes to the gradient. Applying the SDP gradient formula to the SDP section II, where the corresponding Lagrangian is $\mathcal{L}(\rho,\delta)=\delta+\mathrm{Tr}(\rho(H^{2}-\delta H))$, and the primal-dual optimum is ($\rho^{\star}=|\psi_{1}\rangle\langle\psi_{1}|/\Delta,\delta^{\star}=\Delta$), we get
[…]
where $\psi_{1}$ is the first excited state with energy $\Delta$.
We can use the same gradient formula for the LTI SDP in eq. 11. This can be useful for example if we are given a parent Hamiltonian with interaction term $h_{0}$ and we wish to find a different Hamiltonian with the same ground state but with a bigger gap. We can use the solution of the SDP to perform a gradient search over the interaction term $h$ to maximize the gap (while maintaining frustration freeness).
Let $p_{0}$ be the projector to the ground space of $h_{0}$. Any interaction term of the form $h(X):=h_{0}+(\mathbb{I}-p_{0})X(\mathbb{I}-p_{0})$ which is still positive semidefinite and frustration free will have the same global ground state. We can use the SDP gradient formula to differentiate the gap which the SDP detects with respect to $X$. Of course we can always increase the gap arbitrarily by choosing $X=\mathrm{const.}\times\mathbb{I}$, so in addition to the positivity constraint $h(X)\geq 0$ we should impose some constraint on $X$, e.g. $||X||^{2}_{2}\leq 1$. Both of those constraints are easily differentiable and so we could perform constrained gradient optimization.
Let us now consider the case of the LTI gap detection SDP eq. 11. Applying the SDP gradient formula ([39], Section 4.2) we see that the gradient of the gap is given by
[…]
where $g_{n}(X,\delta)=-\delta h_{1}(X)+h_{1}(X)^{2}+\sum_{j=2}^{n-1}\{h_{1}(X),h_{j}(X)\}$.
Appendix E Certifiable formulation of the LTI SDP
We show the steps to ensure that lower bounds produced by the SDP do not suffer from numerical precision errors. For the LTI SDP,
$\Delta_{\mathrm{LTI}}(n)$ $=\max_{\delta,Y}\;\delta$ (133)
s.t. $\quad g_{n}(\delta)+\mathbb{I}\otimes Y-Y\otimes\mathbb{I}\succeq 0.$ (134)
$\quad g_{n,i}(\delta):=\left(h_{i}^{2}+\sum\limits_{i<j<i+n-k}\{h_{i},h_{j}\}-\delta h_{i}\right),$ (135)
the solver produces a solution $\delta$ for which the constraint is positive semidefinite only to some finite precision:
$q_{n}(\delta):=g_{n}(\delta)+\mathbb{I}\otimes Y-Y\otimes\mathbb{I}\succeq-\epsilon,$ (136)
for some small $\epsilon$ (for MOSEK, which is the solver we use, $\epsilon\approx 1e-8$). Such a solution cannot be used to prove a gap in the thermodynamic limit (or for large finite systems) because for a slightly negative $q_{n}$ as in eq. 136, we get
$\sum^{N}_{i}\tau_{i}q_{n}\succeq-N\epsilon$ (137)
. Therefore, to produce rigorous certificates for out of finite-precision solutions, we have to ensure that $q_{n}(\delta)$ is strictly positive semidefinite.
We can make such a certificate if the ground state of $H$ is given to us as a matrix-product-state. Note that for a frustration-free Hamiltonian the ground state is always an MPS if the ground state degeneracy remains bounded for all $N$ (theorem IV.17 in [48]). Moreover, in all practical use-case scenarios we have an MPS description of the ground state. We therefore assume that the MPS tensor describing the ground state is given to us.
To make our solution certifiable we slightly modify the LTI SDP. We proceed as follows: First we note that all the terms in $g_{n}$ annihilate the ground state MPS as the latter is annihilated by each individual term $h_{i}$. Note that, for $q_{n}$ to be a valid generating term for $[H^{2}]_{n}-\delta H$, the extra term $\mathbb{I}\otimes Y-Y\otimes\mathbb{I}$ also has to annihilate the ground state MPS. To ensure this, we construct the projector to the orthogonal complement of the $n-1$ site MPS, which we denote as $P_{n-1}^{\perp}$ (which is also the $(n-1)$-body parent Hamiltonian projector) and replace $Y$ with $P_{n-1}^{\perp}YP_{n-1}^{\perp}$ in eq. 133. This ensures that the whole operator $q_{n}$ annihilates the ground state MPS. The resulting $\delta$ might now be lower because we restricted the set of allowed $Y$s, but this is acceptable, since it still represents a lower bound on the gap. The new form for the certification LTI SDP is then
$\Delta^{\textbf{cert}}_{\mathrm{LTI}}(n)$ $=\max_{\delta,Y}\;\delta$ (138)
s.t. $\quad g_{n}(\delta)+\mathbb{I}\otimes P_{n-1}^{\perp}YP_{n-1}^{\perp}-P_{n-1}^{\perp}YP_{n-1}^{\perp}\otimes\mathbb{I}\succeq 0.$ (139)
$\quad g_{n,i}(\delta):=\left(h_{i}^{2}+\sum\limits_{i<j<i+n-k}\{h_{i},h_{j}\}-\delta h_{i}\right)$ (140)
With the above constraint, we have ensured that the new operator $q_{n}(\delta)$ annihilates the ground state MPS. This means, by construction $q_{n}(\mathbb{I}-P_{n}^{\perp})=0$ as the support of the projector $(\mathbb{I}-P_{n}^{\perp})$ is spanned by the $n$-body MPS (with all possible boundary states). Additionally, to fully ensure that $q_{n}(\delta)\succeq 0$ after the optimization is performed, we check the minimal eigenvalue of $\lambda_{\mathrm{min}}(P_{n}^{\perp}q_{n}(\delta)P_{n}^{\perp})$, which is $q_{n}$ projected to the orthogonal complement of the $n$-body MPS. If $\lambda_{\mathrm{min}}(P_{n}^{\perp}q_{n}(\delta)P_{n}^{\perp})$ is positive, then we have rigorously obtained a lower bound because the $q_{n}$ is zero in on $\mathbb{I}-P_{n}^{\perp}$ and strictly positive on $P_{n}^{\perp}$. On the other hand, if its smallest eigenvalue on $P_{n}^{\perp}$ is negative (we expect it to be small in magnitude because the constraint $q_{n}(\delta)$ can only be violated up to the solver precision) we keep reducing the gap ($\delta$) by a small number, until it results in a strictly positive smallest eigenvalue. If this is the case, we simply take the smaller value as our certified gap. In practice it is usually sufficient to introduce a correction on the order of $\epsilon/\gamma_{2}(h_{i})$, where $\gamma_{2}(h_{i})$ is the gap of the interaction term, to get a strictly positive $q_{n}$.